# concert_supp

### v2

@@ 4.2 Adaptive theoretical results with unknown $T^{*}$
Theorem 6 (Prediction).
Under the assumptions in Theorem 4 and Assumption LABEL:assump:eigen-LABEL:assump:pred in the supplementary material, the posterior satisfies that,
$\Pi\left(\boldsymbol{\beta}^{(0)}:\|X^{(0)}(\boldsymbol{\beta}^{(0)}-\boldsymbol{\beta}^{(0)*})\|_{2}\geq M_{1}\sigma^{(0)}_{y}s^{*}\sqrt{\frac{n_{0}\log p}{n_{S^{*}}}}~{}\bigg{|}~{}\{y^{(k)}\}_{k=0}^{K}\right)\leq\tilde{c}_{3}p^{-1},$
for some constant $\tilde{c}_{3},M_{1}>0$.
Theorem 5 and 6 show that the CONCERT guarantees preferable posterior contraction rates in terms of $\ell_{1}/\ell_{2}$ loss and prediction error under the unknown $T^{*}$ case, which extends the results in Theorem 2 and 3.
@@ 5 Simulation Studies
In this section, we empirically evaluate the performance of our proposed method CONCERT and compare it with various baseline target methods and state-of-art transfer learning methods in extensive simulation studies. Specifically, target methods include Lasso (Tibshirani, 1996), SparseVB with Laplace slab prior (Ray and Szabó, 2022), and NaiveVB where CONCERT simplifies to a solo of target data. Transfer methods include TransLasso (Li et al., 2022) and TransGLM (Tian and Feng, 2023).
We consider both linear and logistic regression and present the estimation errors in three different settings. In Section 5.1, we examine the performance of all methods in the framework of informative set, a scenario commonly adopted in the transfer learning literature. In Section 5.2, we consider the case where sources are heterogeneous and only share partial information with the target. In Section 5.3, all sources share the whole information with the target but with redundant signals.
We also compare the prediction errors and variable selection results in Section LABEL:append:prediction and LABEL:append:selection in Supplementary Material, respectively. The posterior similarity probabilities of CONCERT are investigated in Section LABEL:append:source_var. We study the sensitivity of our CONCERT to hyperparameters and compare the run time with various methods (including the MCMC alternative) in Section LABEL:append:hyper and LABEL:append:time, respectively. Simulations under other settings are conducted in Section LABEL:append:simu.
@@ 5.1 Existence of Informative Set
Motivated by Li et al. (2022), we first consider the scenario where there exists a so-called informative set $\cal A$. If a source $k\in\cal A$, the similarity between this source and the target is high in the sense that the contrast $\boldsymbol{\delta}^{(k)}$ enjoys a small norm and sparse structure. In contrast, if a source $k\not\in{\cal A}$, i.e., is non-informative, its parameter deviates from the target parameter more severely. Note that such non-informativeness is based on a global consideration, while locally, potential covariates in the source data may be helpful.
@@
where $H_{k}$ is a random index subset of $[p]$ with $|H_{k}|=h$ and $\operatorname{Sign}(j)$ takes ±$\pm 1$ with equal probability. If $k\not\in{\cal A}$, we set $\beta^{(k)}_{j}=\beta_{j}^{(0)}+\operatorname{Sign}(j)$ if $j\in H_{k}$ with $|H_{k}|=20$. We consider varying $|{\cal A}|$ and $h$ with $|{\cal A}|\in\{0,2,4,6,8,10\}$ and $h\in\{4,8,12\}$. We run the experiment on 100 independent trials and summarize the means and standard errors of the rooted sum of squared estimation errors of $\boldsymbol{\beta}^{(0)}$ for each method. The results are reported in Figure 2.
@@
As we observe from Figure 2, CONCERT exhibits a significant advantage over all other candidates. We underscore three highlights. First, our CONCERT significantly improves the performance over traditional VB estimators, indicating the benefits of transfer learning. Though target VB methods perform worse than Lasso in the logistic setting, CONCERT outperforms TransGLM considerably. Second, all the transfer learning methods perform worse as the size of informative sources decreases. However, CONCERT is much less sensitive to the decrease of globally informative sources, especially in the linear setting. When all the sources are globally non-informative, CONCERT maintains excellent performance, while other transfer learning methods deteriorate to target Lasso or even show negative transfer. This is because our covariate-specific informative indicators $I_{j}^{(k)}$ can adaptively include locally informative covariates in different sources. Third, CONCERT is also robust to the varying size $h$, which is also due to our partial similarity modeling.
@@ 5.2 Heterogeneous Information
In this subsection we consider a more general scenario where sources contain heterogeneous information and only share some common signals with the target. In the sense of informative set, they should be considered non-informative.
Here the covariates and noises are generated in the same manner as that in Section 5.1. For the parameters, we set $\boldsymbol{\beta}^{(0)}=(\mathbf{0.5}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for linear regression and $\boldsymbol{\beta}^{(0)}=(\mathbf{1}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for logistic regression with varying $s\in\{4,8,12,16\}$, and each $\boldsymbol{\beta}^{(k)}$ only shares partial values of $\boldsymbol{\beta}^{(0)}$. Specifically, we define the informative signal ratio $w$ as the ratio of shared target signals between $\boldsymbol{\beta}^{(k)}$ and $\boldsymbol{\beta}^{(0)}$ among the first $s$ components. Then we set
@@
where $W_{k}$ is a random subset of $[s]$ with $|W_{k}|=\lceil ws\rceil$ and $U_{k}$ is a random subset of $\{s+1,\ldots,p\}$ with $|U_{k}|=20$. We set the ratio to take values in $\{0.6,0.7,0.8,0.9,1\}$ and present the results based on 100 replications in Figures 3.
@@
From Figure 3, it can be seen that the challenge of information transfer intensifies as the target information becomes more complicated and the source heterogeneity increases. However, CONCERT can still improve much compared with NaiveVB and SparseVB. We highlight two points here. Firstly, CONCERT outperforms other transfer methods significantly. As in most cases, all sources are globally non-informative, TransLasso and TransGLM fail to extract useful knowledge from source data, which makes them fail to compete with Lasso on target data. However, our CONCERT succeeds in identifying locally informative covariates in most cases and shows strong robustness when the amount of shared information is not dominant. Secondly, though the performance of all transfer methods improves as the informative signal ratio becomes larger, CONCERT is much more responsive. We also note that the advantage of CONCERT over other candidates becomes much clearer as the sparsity number $s$ increases.
@@ 5.3 Redundant Information
We further investigate the power of CONCERT on information transfer under the scenario where all sources contain complete target signals, however, with various degrees of contamination on the rest. Though intrinsically these sources are all excellent candidates for transfer learning, many of them are regarded as globally non-informative and can be removed under the framework of informative set.
@@
where $\rho$ represents the redundant signal strength with $\rho\in\{0.3,0.5,1,1.5\}$ and $|U_{k}|\in\{4,8,12,16,20\}$ is the redundant signal number. We repeat the experiment 100 times and the summarized results are shown in Figure 4.
@@
Figure 4 shows that CONCERT can avoid being misled by redundant signals and performs the best. While other transfer methods deteriorate rapidly as the redundant signal numbers and strengths grow, CONCERT can identify locally informative covariates, which reduces its bias even when the magnitude of redundant signals dominates that of true signals. We want to highlight that other transfer methods may suffer from negative transfer, especially in the cases of large redundant signal strengths, while CONCERT remains robust in all cases and is generally not affected by redundant signals in linear regression.
@@ 6 Real Data Analysis
In this section, we apply our CONCERT to GTEx data and Lending Club data to demonstrate its effectiveness under the linear and logistic regression settings, respectively.
@@ 6.1 Application to GTEx Data
The proposed transfer learning algorithm is applied to the Genotype-Tissue Expression (GTEx) data, accessible at https://gtexportal.org/. This dataset encompasses gene expression levels from 49 tissues, involving 838 individuals and aggregating 1,207,976 observations of 38,187 genes. Building upon the methodology outlined by Li et al. (2022), our investigation focuses on gene regulations within the central nervous system (CNS) across diverse tissues. The collection of CNS-related genes is consolidated under MODULE_137, comprising a total of 545 genes, alongside an additional 1,632 genes that exhibit significant enrichment in the same experiments as those within the module.
Our specific interest lies in the prediction of the expression level of the gene SH2D2A (SH2 Domain Containing 2A) in a target tissue through the utilization of other central nervous system genes. SH2D2A encodes a T-cell-specific adapter protein that regulates early T-cell activation. It plays a crucial role in inflammatory neuropathies, particularly chronic inflammatory demyelinating polyneuropathy (Uncini et al., 2011). Six brain tissues are considered target tissues and, for each target, the remaining 48 tissues serve as sources. The corresponding models are individually estimated. We first screen the genes by examining their correlation with the expression level of SH2D2A, leaving 149 genes whose correlation coefficients are greater than 0.25. All the expression levels are standardized. The average sample size for the target dataset is 194, while for the sources, it is 17,135. We randomly split the data into a training set (80% of all samples) and a testing set (20%) 50 times, and compare the performance of CONCERT with the same candidates as listed in Section 5. The average prediction error on the testing sets is reported in Table 1.
@@
In general, our CONCERT performs the best among all the methods. We underscore three findings. Firstly, in all studies CONCERT makes improvements over SparseVB and NaiveVB which solely use target information, indicating that at least one of the sources is partially informative. Secondly, CONCERT outperforms all other transfer methods and enjoys a lower standard deviation. Specifically, for the Hippocampus tissue, CONCERT reduces the prediction error of Lasso by more than 20%, whereas TransLasso and TransGLM only reduce it by less than 10%. This suggests that the source datasets for Hippocampus may have limited useful global information, but there are many local pieces of information that are valuable. Such partial information can be captured by CONCERT, thereby further reducing prediction errors.
Lastly, we observe negative transfer in the Nigra tissue when using TransGLM. This may be due to that the source detection method in TransGLM is based on the global similarity gap assumption, which may not hold here. In contrast, the CONCERT remains robust and provides an improvement over target-only methods.
@@ 6.2 Application to Lending Club Data
In this section, we study the issuing of loan defaults using the Lending Club data available through Kaggle. Lending Club is a financial technology company which offers peer-to-peer (P2P) lending services through its online platform, and the Lending Club dataset contains features on both borrowers and loans issued. We consider a binary response based on the loan status, where loans that are not fully paid are denoted as class 1, and otherwise denoted as class 0. Our goal is to predict the risk of loan default within the target domain using information from other source domains.
We focus on the period from 2007 to 2011. After excluding the variables whose values are missing for most records and adding the pairwise interaction terms between the remaining variables, we obtain 3,458 loan records and 190 predictors in six domains: vacation, medical, wedding, education, moving, and house. We consider all domains as our targets and estimate the corresponding models individually. For each target, all other domains are utilized as sources. Similarly to Section 6.1, we randomly split the data into a training set and a testing set 100 times, and compare the CONCERT with the same candidates for the logistic setting in Section 5. The cutoff for binary classification is set as $0.5$, and the average prediction error on the testing sets is reported in Table 2.
@@
Table 2 shows the advantage of the CONCERT on improving the prediction accuracy of loan default on the target domain. It is noteworthy that Bayesian methods such as SparseVB and NaiveVB can perform much better than Lasso when only the target dataset is used. Although TransGLM shows improvements over Lasso, its performance is only comparable to that of the target-only Bayesian methods. In contrast, our method can further enhance prediction accuracy of the target-only Bayesian methods. Specifically, in the domains of vacation and house, the CONCERT further reduces the prediction errors of SparseVB and NaiveVB by about 10%.
@@ 7 Conclusion
In this paper, we proposed CONCERT, a novel Bayesian method to allow partial information transfer. By imposing a conditional spike-and-slab prior together with the traditional spike-and-slab prior on the joint distribution of parameters in the sources and the target, we can achieve variable selection and information transfer simultaneously, robustly, and adaptively. Our theoretical results demonstrate the validity of CONCERT for robust partial information transfer. We make use of the variational Bayes framework to make our algorithm scalable. Numerical experiments and real data analysis show that the CONCERT is robust to source heterogeneity and can better utilize partially informative covariates from the sources to improve target learning.
There are several potential directions for future work. First, although our simulation results demonstrate that our method performs well in both linear regression and logistic regression, the current theoretical framework is limited to linear regression. Extending the theoretical properties to the generalized linear models might require additional efforts. Second, inference procedures may be developed, which is a significant advantage of Bayesian methods. Moreover, the innovative use of the conditional spike-and-slab prior is not limited to transfer learning. This idea can also be applied to other settings where both shared components and subject-specific components exist, such as data integration or robust learning. These extensions are beyond the scope of this paper and require further development.

### v3

@@ 4.2 VB Posterior Contraction
The results in Theorem 5 include an extra term of $O(1/\kappa_{n})$ in the contraction rates and an extra factor of $\kappa_{n}$ in the upper bounds, compared to the true posterior counterpart in Theorem 3. Since $\kappa_{n}$ diverges at an arbitrarily slow rate, the VB estimator converges only slightly more slowly. Similar insights have been provided in ray2020vb and ray2022vb for single-task Bayesian regression.
@@ 5 Simulation Studies
In this section, we empirically evaluate the performance of our proposed method, CONCERT, and compare it with various baseline target methods and state-of-the-art transfer learning methods in extensive simulation studies. Specifically, target methods include Lasso (tibshirani1996), SparseVB with Laplace slab prior (ray2022vb), and NaiveVB where CONCERT simplifies to a solo of target data. Transfer learning methods include TransLasso (li2022trans) and TransGLM (tian2022transfer).
We consider both linear regression and logistic regression and present the estimation errors in three different settings. In Section 5.1, we examine the performance of all methods in the framework of informative set, a scenario commonly adopted in the transfer learning literature (li2022trans). In Section 5.2, we consider the case where sources are heterogeneous and only share partial information with the target. In Section 5.3, all sources share the whole information with the target but with redundant signals.
Additional simulation results are provided in the Supplementary Material. To demonstrate the robustness of our method, we present a sensitivity analysis of the spike-and-slab hyperparameters by exploring various reasonable combinations of $(q_{0},\eta)$ and $(q_{k},\tau_{k})$ in Section D.4, where our CONCERT consistently achieves satisfactory estimation errors. We discuss the alternative uses of hyperpriors and compare the accuracy and scalability of our VB approach with MCMC alternatives in Section D.5 and D.6, respectively. The R implementation of our method is publicly available at https://github.com/RuqianZhang/CONCERT.
@@ 5.1 Existence of Informative Set
Motivated by li2022trans, we first consider the scenario where there exists a so-called informative set $\cal A$. If a source $k\in\cal A$, the similarity between this source and the target is high in the sense that the contrast $\boldsymbol{\delta}^{(k)}$ enjoys a small norm and sparse structure. In contrast, if a source $k\not\in{\cal A}$, i.e., is non-informative, its parameter deviates from the target parameter more severely.
@@
where $H_{k}$ is a random index subset of $[p]$ with $|H_{k}|=h$ and $\operatorname{Sign}(j)$ takes $\pm 1$ with equal probability. If $k\not\in{\cal A}$, we set $\beta^{(k)}_{j}=\beta_{j}^{(0)}+\operatorname{Sign}(j)$ if $j\in H_{k}$ with $|H_{k}|=20$. We consider varying $|{\cal A}|$ and $h$ with $|{\cal A}|\in\{0,2,4,6,8,10\}$ and $h\in\{4,8,12\}$. We run the experiment on 100 independent trials and summarize the means and standard errors of the root sum of squared estimation errors of $\boldsymbol{\beta}^{(0)}$ for each method. The results are reported in Figure 2.
@@
As we observe from Figure 2, CONCERT exhibits a significant advantage over all other candidates. We underscore three highlights. First, our CONCERT significantly improves the performance over single-task VB estimators, indicating the benefits of transfer learning. Though NaiveVB methods perform worse than Lasso in the logistic setting, CONCERT outperforms TransGLM considerably. Second, all the transfer learning methods perform worse as the size of globally informative sources decreases. However, CONCERT is much less sensitive to the decrease of $|\cal{A}|$, especially in the linear setting. When all the sources are globally non-informative, CONCERT maintains excellent performance, while other transfer learning methods deteriorate to the single-task Lasso, or even show negative transfer. This is because the covariate-specific informative indicators $I_{j}^{(k)}$ in CONCERT can adaptively include locally informative covariates in different sources. Third, CONCERT is robust to the varying size $h$. As $h$ grows, other transfer learning methods show larger estimation errors, especially when $|\cal{A}|$ is large. In contrast, the CONCERT is generally unaffected by $h$ since our partial similarity modeling is able to overcome the deterioration of the globally informative sources.
@@ 5.2 Heterogeneous Information
In this subsection, we consider a more general scenario where sources contain heterogeneous information and only share some common signals with the target. In the sense of informative set in li2022trans, this scenario is considered non-informative.
The covariates and noises are generated in the same manner as in Section 5.1. For the parameters, we set $\boldsymbol{\beta}^{(0)}=(\mathbf{0.5}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for linear regression and $\boldsymbol{\beta}^{(0)}=(\mathbf{1}^{\top}_{s},\mathbf{0}^{\top}_{p-s})^{\top}$ for logistic regression with varying $s\in\{4,8,12,16\}$, and each $\boldsymbol{\beta}^{(k)}$ only shares partial values of $\boldsymbol{\beta}^{(0)}$. Specifically, we define the informative signal ratio $w$ as the ratio of shared target signals between $\boldsymbol{\beta}^{(k)}$ and $\boldsymbol{\beta}^{(0)}$ among the first $s$ components. Then we set
@@
where $W_{k}$ is a random subset of $[s]$ with $|W_{k}|=\lceil ws\rceil$ and $U_{k}$ is a random subset of $\{s+1,\ldots,p\}$ with $|U_{k}|=20$. We set $w$ to take values in $\{0.6,0.7,0.8,0.9,1\}$ and present the results based on 100 replications in Figures 3.
@@
From Figure 3, it can be seen that the challenge of information transfer intensifies as the target information becomes more complicated and the source heterogeneity increases. However, CONCERT demonstrates significant improvement compared to NaiveVB and SparseVB. We highlight two points here. Firstly, CONCERT outperforms other transfer learning methods significantly. As all sources are globally non-informative in most cases, TransLasso and TransGLM fail to extract useful knowledge from source data, which makes them fail to compete with Lasso on target data. In contrast, CONCERT succeeds in identifying locally informative covariates in most cases and shows strong robustness when the amount of shared information is not dominant. Secondly, though the performance of all transfer learning methods improves as the informative signal ratio becomes larger, CONCERT is much more responsive. Additionally, we note that the advantage of CONCERT over other candidates becomes more apparent as the sparsity number $s$ increases.
@@ 5.3 Redundant Information
We further investigate the scenario where all sources contain complete target signals, however, with various degrees of contamination on the rest. Though intrinsically these sources are all excellent candidates for transfer learning, many of them are regarded as globally non-informative and will be removed under the framework of informative set.
@@
where $\rho$ represents the redundant signal strength with $\rho\in\{0.3,0.5,1,1.5\}$ and $|U_{k}|\in\{4,8,12,16,20\}$ is the redundant signal number. We repeat the experiment 100 times, and the summarized results are shown in Figure 4.
@@
Figure 4 shows that CONCERT can avoid being misled by redundant signals and performs the best. While other transfer learning methods deteriorate rapidly as the redundant signal numbers and strengths grow, CONCERT can identify locally informative covariates, which reduces its bias even when the magnitude of redundant signals dominates that of true signals. We highlight that other transfer learning methods may suffer from negative transfer, especially in the cases of large redundant signal strengths, while CONCERT remains robust in all cases and is generally not affected by redundant signals in linear regression.
@@ 6 Real Data Analysis
In this section, we apply CONCERT to GTEx data and Lending Club data to demonstrate its effectiveness under the linear and logistic regression settings, respectively.
@@ 6.1 Application to GTEx Data
The proposed transfer learning algorithm is applied to the Genotype-Tissue Expression (GTEx) data, accessible at https://gtexportal.org/. This dataset encompasses gene expression levels from 49 tissues, involving 838 individuals and aggregating 1,207,976 observations of 38,187 genes. Building upon the methodology outlined by li2022trans, our investigation focuses on gene regulations within the central nervous system (CNS) across diverse tissues. The collection of CNS-related genes is consolidated under MODULE_137, comprising a total of 545 genes, alongside an additional 1,632 genes that exhibit significant enrichment in the same experiments as those within the module.
Our specific interest lies in the prediction of the expression level of the gene SH2D2A (SH2 Domain Containing 2A) in a target tissue through the utilization of other central nervous system genes. SH2D2A encodes a T-cell-specific adapter protein that regulates early T-cell activation. It plays a crucial role in inflammatory neuropathies, particularly chronic inflammatory demyelinating polyneuropathy (Uncini2011polumorphism). Six brain tissues are considered as target tissues and, for each target, the remaining 48 tissues serve as sources. The corresponding models are individually estimated. We first screen the genes by examining their correlation with the expression level of SH2D2A, leaving 149 genes whose correlation coefficients are greater than 0.25. All the expression levels are standardized. The average sample size for the target dataset is 194, while for the sources, it is 17,135. We randomly split the data into a training set (80% of all samples) and a testing set (20%) 50 times, and compare the performance of CONCERT with the same candidates as listed in Section 5. The average prediction errors on the testing sets are reported in Table 1.
@@
In general, CONCERT performs the best among all the methods. We underscore three findings. Firstly, in all studies, CONCERT makes improvements over SparseVB and NaiveVB which solely use target information, indicating that at least one of the sources is partially informative. Secondly, CONCERT outperforms all other transfer learning methods and enjoys a lower standard deviation. Specifically, for the Hippocampus tissue, CONCERT reduces the prediction error of Lasso by more than 20%, whereas TransLasso and TransGLM only reduce it by less than 10%. This suggests that the source datasets for Hippocampus may have limited useful global information, but there are many local pieces of information that are valuable. Such partial information can be captured by CONCERT, thereby further reducing prediction errors.
Lastly, we observe negative transfer in the Nigra tissue when using TransGLM. This may be due to the fact that the source detection method in TransGLM is based on the global similarity gap assumption, which may not hold here. In contrast, the CONCERT remains robust and provides an improvement over target-only methods.
@@ 6.2 Application to Lending Club Data
In this subsection, we study the issuing of loan defaults using the Lending Club data available through Kaggle at https://www.kaggle.com/datasets/wordsforthewise/lending-club/. Lending Club is a financial technology company which offers peer-to-peer (P2P) lending services through its online platform, and the Lending Club dataset contains features on both borrowers and loans issued. We consider a binary response based on the loan status, where loans that are not fully paid are denoted as class 1, and otherwise denoted as class 0. Our goal is to predict the risk of loan default within the target domain using information from other source domains.
We focus on the period from 2007 to 2011. After excluding the variables whose values are missing for most records and adding the pairwise interaction terms between the remaining variables, we obtain 3,458 loan records and 190 predictors in six domains: vacation, medical, wedding, education, moving, and house. We consider all domains as our targets and estimate the corresponding models individually. For each target, all other domains are utilized as sources. Similarly to Section 6.1, we randomly split the data into a training set and a testing set for 50 times, and compare the CONCERT with the same candidates for the logistic setting in Section 5. The cutoff for binary classification is set as $0.5$, and the average prediction errors on the testing sets are reported in Table 2.
@@
Table 2 shows the advantage of the CONCERT on improving the prediction accuracy of loan default on the target domain. It is noteworthy that Bayesian methods such as SparseVB and NaiveVB can perform much better than Lasso when only the target dataset is used. Although TransGLM shows improvements over Lasso, its performance is only comparable to that of the single-task Bayesian methods. In contrast, our method can further enhance the prediction accuracy of the single-task Bayesian methods. Specifically, in the domains of vacation and house, the CONCERT reduces the prediction errors of SparseVB and NaiveVB by about 10%.
@@ 7 Conclusion
In this paper, we propose a novel Bayesian method to allow partial information transfer. By imposing a conditional spike-and-slab prior together with the traditional spike-and-slab prior on the joint distribution of parameters in the sources and the target, we can achieve variable selection and information transfer simultaneously, robustly, and adaptively. Our theoretical results demonstrate the validity of CONCERT for robust partial information transfer. We adopt the variational Bayes framework to ensure scalability of the algorithm. Numerical experiments and real data analysis show that the CONCERT is robust to source heterogeneity and can better utilize partially informative sources to improve target learning.
There are several potential directions for future work. First, although the simulation results show strong performance in both linear and logistic regression models, our current theoretical framework only covers the linear case. Extending the theory to logistic regression requires handling the parameter-dependent Hessian that arises from the nonlinearity of the model, which can be addressed using techniques in Narisetty2019ss. Second, inference procedures may be developed, which is a significant advantage of Bayesian methods. Moreover, the innovative use of the conditional spike-and-slab prior is not limited to transfer learning. This idea can be further applied to other settings where both shared components and subject-specific components exist, such as data integration or robust learning. These extensions are beyond the scope of this paper and require further development.
Acknowledgements:
The authors would like to thank the reviewers, an Associate Editor and the Editor for their constructive comments that improved the quality of this paper.
Disclosure statement:
The authors report there are no competing interests to declare.